The probability of passing exam is 0.8, what is the probability of failing exam?
Answer by ikleyn(52879)

.

The probability of failing exam is the COMPLEMENT to the probability of passing it.


        P(failing) = 1 - P(passing) = 1 - 0.8 = 0.2.    ANSWER

Solved.


==================


To better memorize,  keep in mind this joking statement:

        If the probability for a sandwich to fall butter side up is  0.2,
        then the probability for the sandwich to fall butter side down
        is the  COMPLEMENT  to it.